OpenStudy (anonymous):

simplify 3x(4x^4 – 5x)

OpenStudy (anonymous):

12x5−15x2

OpenStudy (anonymous):

Sorry its direct, I really have to go tho

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \color{red}{\rm a}(b-c)=\color{red}{\rm a}\cdot b-\color{red}{\rm a}\cdot c }\) \(\large\color{black}{ \displaystyle x^{\rm m} \cdot x^{\rm n}=x^{\rm m+n}}\)

OpenStudy (tanya123):

12 x^5-15 x^2 x^2 (12 x^3-15)

OpenStudy (anonymous):

thank you :)